[Hallc_running] RC Update

Dipangkar Dutta ddutta07 at gmail.com
Sun Jan 21 23:03:08 EST 2018

Hi All,
The LH2 target in Hall C is being warmed up right now, so that Dave Meekins
can try to replace the 6% carbon foil on the solid target ladder. The
scattering chamber will be rad surveyed tomorrow morning and then opened
up. The work is expected to take a minimum of 5 shifts. The owl shift
tonight is cancelled and the day shift tomorrow will need one person in the
counting house. More information about the shift schedule for the next few
days will be sent tomorrow by the new RC ( Latif ).

The F2 experiment collected some data using the HMS during the swing shift
and some of the misc. SHMS commissioning tasks were completed too.

There will be no RC meeting tomorrow because of the Hall C collaboration

(Hall C RC)
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180121/bba332fb/attachment.html>

More information about the Hallc_running mailing list